# Explainable AI for Autism Diagnosis: Identifying Critical Brain Regions Using fMRI Data

### v1

@@ V Conclusion
Our feature analysis highlighted the importance of visual processing regions, particularly the calcarine sulcus and cuneus, in ASD classification. These findings were consistent across different preprocessing pipelines and aligned with independent genetic studies [68], providing strong validation of our model’s ability to capture genuine neurobiological markers of ASD rather than dataset-specific patterns. Notably, our identified biomarkers were common among a variety of ASD severity cases in our dataset, suggesting these visual processing regions may represent a fundamental neurobiological signature of ASD across the spectrum of severity.
In summary, this research has made significant strides in the application of deep learning to ASD classification and biomarker identification. By developing an accurate, explainable model and rigorously benchmarking interpretability methods, we have not only advanced the technical aspects of ASD diagnosis but also contributed to the neurobiological understanding of the disorder. The consistent identification of visual processing regions as key discriminators across different severity levels of ASD opens new avenues for research and potentially for clinical applications. While challenges remain, this work represents a substantial step forward in the quest for objective, reliable, and interpretable methods for ASD diagnosis and characterisation.
@@ VI Limitations and Future Directions
Despite the promising results of this study, several limitations must be acknowledged, each of which points to important directions for future research. A primary limitation is the binary nature of our classification model. By focusing on distinguishing between ASD and TC and due to a dataset limitation, we may have inadvertently created a classifier that identifies general neurodivergence rather than ASD specifically. This raises the question of whether our model can truly differentiate ASD from other neurodevelopmental disorders or atypical brain patterns. Future research should explore multi-class classification models that can distinguish ASD not only from typical controls but also from other neurodevelopmental conditions.
Secondly, while our model achieved high accuracy and identified consistent biomarkers, the dataset used was limited in the amount of information given, particularly in terms of ethnic representation. This limitation potentially affects the generalisability of our findings to the broader ASD population and its potential to perpetuate healthcare disparities [75]. Future studies should prioritise the use of explicitly diverse and representative datasets to ensure that the identified biomarkers and classification methods are valid across different demographic groups.
An important future direction would be to develop a model that can highlight important regions for ASD classification on a per-MRI scan basis, rather than studying the model’s overall behavior. This approach could provide more personalised insights into the neurobiological markers of ASD for each individual, potentially revealing how these markers vary across the spectrum and among different demographic groups. This could potentially also aid clinicians with early and accurate diagnosis of ASD.
Furthermore, a multimodal approach holds significant promise. Integrating behavioural, cognitive, genetic, and structural imaging data alongside functional connectivity could provide a more holistic and nuanced picture of ASD [76]. This could lead to the identification of ASD subtypes or personalised profiles, enabling tailored interventions.
Finally, collaborative research models are needed to bridge the gap between AI development and clinical practice. Interdisciplinary teams of neuroscientists, AI researchers, and clinicians are crucial to design studies with translational impact, ensuring that research questions are clinically relevant and AI models are designed to address the complexities of real-world ASD diagnosis and care.

### v2

@@ V Conclusion
Our feature analysis highlighted the importance of visual processing regions, particularly the visual processing regions (calcarine sulcus and cuneus), in ASD classification. These findings were consistent across different preprocessing pipelines and aligned with independent genetic studies [64], providing strong validation of our model’s ability to capture genuine neurobiological markers of ASD rather than dataset-specific patterns. Notably, our identified biomarkers were common among a variety of ASD severity cases in our dataset, suggesting these visual processing regions may represent a fundamental neurobiological signature of ASD across the spectrum of severity.
In summary, this research has made significant strides in the application of deep learning to ASD classification and biomarker identification. By developing an accurate, explainable model and rigorously benchmarking interpretability methods, we have advanced the technical aspects of ASD diagnosis and contributed to the neurobiological understanding of the disorder. The consistent identification of visual processing regions as key discriminators across different severity levels of ASD opens new avenues for research and potentially for clinical applications. While challenges remain, this work represents a substantial step forward in the quest for objective, reliable, and interpretable methods for ASD diagnosis and characterisation.
@@ VI Limitations and Future Directions
Despite the promising results of this study, several limitations must be acknowledged, each of which points to important directions for future research. A primary limitation is the binary nature of our classification model, which currently only distinguishes ASD from typical controls. By focusing on a binary classification, we risk overlooking the unique characteristics of other neurodevelopmental conditions that could share some of the same features.
Although we report strong results in the ABIDE dataset, testing on other multi-center, ethnically diverse rs-fMRI datasets or prospectively acquired clinical data is essential to confirm real-world portability. We plan to extend our pipeline to these data in future work, especially considering the considerable resource requirements for new fMRI scans. In principle, if new data uses a similar preprocessing pipeline and parcellation scheme, the model would not need retraining. Otherwise, we anticipate that partial or complete retraining will be required.
Another future direction would be to develop a model that can highlight important regions for ASD classification on a per-MRI-scan basis, rather than only analysing overall group differences. This individualised approach could enhance interpretability and identify brain regions that are relevant to a specific patient’s presentation, potentially helping clinical trust and targeted intervention strategies.
Additionally, a multimodal approach holds significant promise. Integrating behavioural, cognitive, genetic, and structural imaging data alongside functional connectivity could provide a more holistic and nuanced picture of ASD. This approach may lead to the identification of ASD subtypes or personalised profiles, enabling more effective interventions. Finally, collaborative research is needed between neuroscientists, AI researchers, and clinicians to design clinically relevant studies and ensure that AI-driven diagnostic methods account for real-world complexities, including data diversity, and ethical deployment of AI in clinical settings.